OpenStudy (anonymous):

Which function below is the inverse of f(x) = The quantity of three x plus five, over seven.?

OpenStudy (anonymous):

y=(3x+5)/7 switch x and y then solve for y, then you'll have the inverse

OpenStudy (anonymous):

ok thank you so much

OpenStudy (anonymous):

x=(3y+5)/7 7x=3y+5 7x-5=3y (7x-5)/3=y